You are here: Foswiki>PartnerObserving Web>TechNightLogs>December20UT (21 Jul 2009, NormCushing)Edit Attach
20 December 2008 UT

Summary

The humidity dropped and skies cleared but there was too much ice on the enclosure to open. In any case, there was much closed-dome testing that could be done and the night was used running OBs which simulated science OBs to test the operation of the Blue and Red cameras - in particular to try to reproduce the Blue contoller errors we had seen the past 3 nights. This was not successful - the Blue camera ran without error all night, till 5am when we decided to call it a night. I am not sure what action taken during the day caused this improvement, but, during the morning and afternoon the Blue Science controller fans were cleaned or revived in some way (it was found that they had stopped working) and the fiber optic cables were secured. Upgrades to the firmware were implemented. Also, the treehouse setpoint temperature was increased to 5C - this was to aid startup after a long periods (~several weeks or a month) when LBC is not used and the PCs are off.

We did investigate a problem of incorrect position angle for the first exposures (red and blue) after a slew - see issue #1462. I believe that the parallactic angle which LBC uses for this exposure is the same as what it used for the last exposure before the slew. I noticed also that the data dictionary values of the target & achieved parallactic angle both vary during the slew, while for comparison, the target DEC value is fixed and only the achieved DEC value varies during the slew.

The network is very slow - editing a simple text file using "vi" can take several seconds to save changes. This is on lbtmu03 and the file is in my account.

PartnerObserving Log

morning/afternoon Andrea worked with SkyTech to get new updates for the controller firmware. Last night, it was not possible to run 'testcamera' --- attemps to take an exposure ended with 'seg fault'. testcamera working now.

~14:30 Going up to LBC-Blue camera to look for any loose fiber from the science controller which may be interfering with communications at >0 deg elevation and range of rotator angles. Wiggling fiber while taking an image with testcamera did not affect the exposure or readout.

afternoon Discussion about how to recover from a communications failure. If it occurs during readout (as opposed to during preset or expose), there is no feedback to the CMU, so the observation cannot be repeated. If it occurs during preset or expose, functions in the CMU, then a failure could trigger a initialization or restart of observation. My sketchy interpretation...

~17:00 This afternoon Fernando noticed that the fans on the Blue Science controller were not running. They replaced these with fans (or some part) from the Red science spare controller. Perhaps the lack of cooling was causing the communication problems?

~20:00 Aaron changing the setpoint temperature of the upper left treehouse to 5C. There was discussion this afternoon about the difficulty starting the LBC computers after a long time off (~1 month). The treehouse can get down to 0C and there is condensation on the boards. In November the LBC team had to use a heat gun to remove condensation and start the PCs.

20:45 Vincenzo sending OB to test communications problems. We noted that the first exposure was done with the rotator angle = 90 and both PA and parallactic angle reported by the User Interface were incorrect. All were changing during the exposure - it looked as though the rotator was moving during the exposure but we do not have the dome open so cannot verify by star trails. By the 2nd exposure, the PA, parallactic angle and rotator angle reported on the User Interface had reached correct position. The image number is lbcb.20081220.034821.fits and the rotator, parallactic and position angle are the following. Note that the parallactic and PA that are in the headers (and below) for 034821 are not the same as what the User Interface reported during the observation.

lbcb.20081220.034821.fits[1]    89.15023        89.00678        -0.14345
lbcb.20081220.034933.fits[1]    -62.52262       -62.52433       -0.00171
lbcb.20081220.035015.fits[1]    -62.27213       -62.30131       -0.02918
lbcb.20081220.035058.fits[1]    -62.13590       -62.16813       -0.03223

Andrea thinks that the TCS is sending the actual and not target value for parallactic angle during the slew and he said that this problem started after he edited the LBC code to move the rotator during the slew, so as to improve efficiency.

21:30 No problems as yet with the Blue controller. So far 3 OBs have completed successfully: one at PA=0 and another at PA=-90. Elevations 70 and 50. One is now sent with PA=180, RA=-255 and elevation=50 degrees.

22:19 problem with red filters - did not position reach position by the maximum of 10 attempts.

00:00 continue running multi-filter, multi-position OBs at a range of elevation and rotator angles - particularly trying to reproduce the position at which the ccd controller went into an error state last night. So far it is working fine. These OBs are in /home/lbcobs/ObservingBlocks/TESTOBS

~2:00 but throughout night, actually Slowness in computer. I am logged into lbtmu03 as okuhn and editing a file on /home/okuhn with vi --- it sometimes takes about 5 seconds to save the file.

03:00 blue camera has worked all night so far for Vincenzo. I'll continue to exercise it to see how it goes. 5-dither, 4-filter binocular OB at elev=87deg and rotator angle=-58 deg. But... note that again on the first exposure, rotator angle = 76, PA = -134, para = -57 ang changing during exposures: 101246(B) and 101244(R).

03:45 moving slightly so as still to be near zenith, but this time use rotator angle = 0. Well, I slewed and stopped to read the parallactic angle and so to set the position angle equal to that to make the rotator angle 0: rotator = parallactic - position. I tried to send -17 = PA, but it did not go there and instead settled at about -13. Tried again - this time sent PA=39 (parallactic angle had changed to about 39) and the position angle for the first 2 exposures is ~42 degrees instead.

04:20 Moved to N part of sky and HA=0. Par=-180, PA=150, Rot=50 or so for the first exposure - incorrect, but for the second (B 112336) it is OK. OB was running well but I stopped it to move on to check the parallactic angle problem for the first integration.

04:40 Slewing telescope and monitoring DDViewer value for achieved parallactic angle under the suspicion that this is what is grabbed, instead of target parallactic angle, for the first integration (can I view 2 data dictionary variables simultaneously?). *Both DDViewer values - for target and achieved parallactic angle were -142, correct. But, the first image was exposed while the rotator=-187 and PA=45 (according to the User Interface). These images are 114112(B) and 114114(R). On the 2nd exposures after the slew (114257 B and R), the PA=0 and rotator angle =-142, correctly.

04:49 Watching target parallactic angle value in DD during slew - it changes and is not a constant value corresponding to the target position (12:00:00, 10:00:00). target=achieved value during the slew. Is this correct? First set of exposures during this sequence (ParAngle_test2) start with parallactic = -39, position= 104 deg and rotator= -143 deg according to User Interface. 115206 (B) and 115204 (R). Before second set (115310 B and 115305 R) rotator angle and position angle were correctly reported as PA=0, rotator=-39 deg.

05:00 Made another long slew. Verified that target DEC is fixed and achieved DEC is changing during slew. Found again that the first set of observations (120308 B and were not at the PA commanded, PA=0, but instead at PA=-95, rotator angle = -37 and parallactic angle = -132. The second set were correctly at PA=0, rotator angle = -132 and parallactic angle = -132. It looks like the parallactic angle assumed for the first image after a slew is the same as it was for the last position from the previous image before the slew. See, e.g. the transition between lbcr.20081220.040456 and 040859 (the angles are similar for the corresponding Blue images):

          image                                                      lbcobnam        lbcobfil                              ra                     dec             rotangle        pa_pnt          parangle
/Repository/lbcr.20081220.040456.fits[1]        allfiltnodither lbc.081220.035909.ob    03:25:00.001    -06:00:00.00       75.70898        -90.02533       -14.31634
/Repository/lbcr.20081220.040859.fits[1]        allfilt3dither3 lbc.081220.040643.ob    06:15:00.000    +32:00:00.00    -193.65381      180.04652       -13.60728
/Repository/lbcr.20081220.041004.fits[1]        allfilt3dither3 lbc.081220.040643.ob    06:14:59.999    +32:00:00.00    -254.47852      180.00562       -74.47289
/Repository/lbcr.20081220.041056.fits[1]        allfilt3dither3 lbc.081220.040643.ob    06:15:00.001    +32:00:00.00    -254.55762      180.00590       -74.55172
/Repository/lbcr.20081220.041149.fits[1]        allfilt3dither3 lbc.081220.040643.ob    06:15:00.001    +32:00:00.00    -254.62354      180.00772       -74.61582

and

/Repository/lbcb.20081220.115836.fits[1]        ParAngle_test2  lbc.081220.114732.ob    11:59:58.646    +10:01:00.00    -37.21012       -0.04925   -37.25937
/Repository/lbcb.20081220.120308.fits[1]        ParAngle_test3  lbc.081220.120024.ob    12:00:00.000    +49:59:59.99    -37.55289       1.01735    -36.53554
/Repository/lbcb.20081220.120409.fits[1]        ParAngle_test3  lbc.081220.120024.ob    12:00:00.000    +50:00:00.00    -132.61979      0.00312 -132.61666
/Repository/lbcb.20081220.120454.fits[1]        ParAngle_test3  lbc.081220.120024.ob    12:00:00.000    +50:00:00.01    -133.14713      0.06593 -133.08120
/Repository/lbcb.20081220.120540.fits[1]        ParAngle_test3  lbc.081220.120024.ob    11:59:55.854    +49:58:39.99    -133.49869      0.06949 -133.42920

There is a strange case where it is in between, but then also the ra and dec are in between. This may be a case where I stopped the OB and the first image was taken before the telescope reached target.
/Repository/lbcb.20081220.044636.fits[1]        allfilt3dither3 lbc.081220.043329.ob    06:15:00.446    +31:59:31.33    -77.05094       0.01573 -77.03521

/Repository/lbcb.20081220.044948.fits[1]        allfilt5dither  lbc.081220.044806.ob    04:19:36.037    +14:06:05.94    -40.01822       -0.00539        -40.02361
/Repository/lbcb.20081220.045148.fits[1]        allfilt5dither  lbc.081220.044806.ob    01:49:59.143    +25:00:24.50    62.49740         0.00408        62.50148
/Repository/lbcb.20081220.045328.fits[1]        allfilt5dither  lbc.081220.044806.ob    01:50:00.000    +25:00:00.00    64.80453        -0.00965        64.79488
/Repository/lbcb.20081220.045508.fits[1]        allfilt5dither  lbc.081220.044806.ob    01:50:00.000    +25:00:00.00    65.00668        -0.01507        64.99161
/Repository/lbcb.20081220.045648.fits[1]        allfilt5dither  lbc.081220.044806.ob    01:50:00.000    +25:00:00.00    65.20004        -0.02063        65.17941

20 December AM Fernando & Vincenzo noted that the Blue Science fans that they 'revived' in the afternoon were spinning more slowly than yesterday and so they are going to replace them with fans from the Red Spare controller.

-- OlgaKuhn - 19 Dec 2008
Topic revision: r7 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback